An impure sample of CaCO3 ( calcium carbonate ) contains 38% of Ca . The percentage of impurity present in the sample is ?

Grade:11

2 Answers

Jatin yadav
6 years ago
100 % pure CaCO3 should contain 40 % of Ca ... But it contains 38 % ..therefore ,Percentage purity = 38/40 x 100 = 95 % ..Impurity = 5 %
Sumedh Meshram
4 years ago
Molecular Weight(CaCO3) - 40+12+48=100gm
Ca%=40÷100 ×100
        =40%
%Impurity=2/100÷40/100×100
                 =2/40×100
                 =5%
The percentage impurity is 5%.